OpenStudy (anonymous):

Simplify. 7 – (–12) – 5 A. 0 B. 24 C. –10 D. 14

OpenStudy (shamil98):

7 -(-12) - 5 - x - = + (7 + 12) - 5 =

OpenStudy (anonymous):

14

OpenStudy (anonymous):

a. 0

OpenStudy (anonymous):

7+12-5=14